OpenStudy (anonymous):

hello! ca somebady help help me , how to do this please -1/3+-7/8 ????

OpenStudy (anonymous):

Hello there. So first of all, do you know what a +- turns into when they are next to each other like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i'm new in this page

OpenStudy (anonymous):

Welcome, by the way. Anyway, so does it go to a positive sign or negative sign?

OpenStudy (anonymous):

im realy bad in this

OpenStudy (anonymous):

Okay thats fine, I just wanted to see if you knew it or not. So when +(-) that will turn the sign into a -

OpenStudy (anonymous):

So now our equation will be -1/3-7/8

OpenStudy (anonymous):

So because both the 1/3 and 7/8 are negative we know that we can almost treat this like an addition problem.

OpenStudy (anonymous):

So when you add 1/3 and 7/8 together you have to first find a common denomiator, correct?

OpenStudy (anonymous):

yes but i cant

OpenStudy (anonymous):

Okay thats fine. One way to find a common denomiator is to multiply the numbers in the denominator together. What will you get when you do that?

OpenStudy (anonymous):

can you do an example with all the instrucion so i can see what you ddi please

OpenStudy (anonymous):

Okie dokie! Good idea!

OpenStudy (anonymous):

Let's say our problem is -5/6+-3/4.

OpenStudy (anonymous):

Since both of these numbers are the same sign, we can add them together as long as we keep the sign. For example if you were to add two positive numbers, you'll get a positive number (example: 3+4=7). If you were to add two negative numbers, you'll get a negative number (example -3+-4=-7). However, if you add a positve number with a negative number you will treat i like a subtraction problem (-13+6=-7, this looks similar to 13-6=7). Anyway, for this particular problem and your problem we are working with two negative signs.

OpenStudy (anonymous):

just became your fan DrPepper, thats for being patient with any type of questions

OpenStudy (anonymous):

thanks for taking the time to help me!!

OpenStudy (anonymous):

Mhm, sorry my computer just froze. :) Anyway, continuing on..

OpenStudy (anonymous):

So now that I know that I will treat this like an addition problem only with a negative sign, I now need to make my fractions have the same number on the bottom of the fraction. So I can multiple whatever is on the bottom of the two fractions together to get this new number. So 6*4=24. So 24 will be on the bottom of my fraction.

OpenStudy (anonymous):

Now since I changed the bottom pparts of the two fractions I now need to change the top parts. Since for 5/6, I multiplied the 6 on the bottom with 4 (as said in the previous paragraph), I need to multiply 4 with the top number 5. So 4*5=20

OpenStudy (anonymous):

Likewise, since for 3/4 I multiplied the 4 on the bottom with a 6 (as said before) I need to multiply the top number in this fraction by 6 as well. So 3*6=18.

OpenStudy (anonymous):

Now my fractions look like this: -18/24+-20/24

OpenStudy (anonymous):

Since the it is 24 on the bottom already I know that my answer will have a 24 on the bottom of the fraction. Now we need to figure out what is on top. So as said this will be like addition. So 18+20=38. However, since both of these numbers are negative we will have a negative sign in front as well, so it will be -38. So combining all of what I said together, we have -38/24.

OpenStudy (anonymous):

which is your unsimplified answer. Does your teacher want your answers to be simplified?

OpenStudy (anonymous):

(@andrius Thank you!)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok. So the last step is to figure out how to simplify this fraction. Well one thing that is noticeable is 38 and 24 are both even numbers, so you can at least divide both 38 and 24 by 2 and see if that will simplify your answer enough.

OpenStudy (anonymous):

So 38/2=19 and 24/2=12. So our fraction is -19/12, and 19 is a prime number (no numbers besides 1 and 19 can divide 19).

OpenStudy (anonymous):

And that is our answer: -19/12.

OpenStudy (anonymous):

Now how about you try it with your problem, step by step, and let me know what you are doing for each step and I'll tell you if you are on track or not. Okay?

OpenStudy (anonymous):

DrPepper I don't know if you are a teacher or not, but sure are a great pedagog!

OpenStudy (anonymous):

\[-1/3+-7/8 \] \[[-(1)(8) - (3)(7)]/ (3)(8)\] \[(-8-21) / 24\] \[-29/24\]

OpenStudy (anonymous):

marxian, funny you should say that. My major in college is actually secondary education mathematics so your comment is very much encouraging. :D